# Solving dynamic portfolio selection problems via score-based diffusion models

### v1

@@ 4.2 Solving portfolio choice problem with $\mathbb{Q}$
For reinforcement learning in financial applications, it is advisable to refrain from utilizing long-term data, as certain global factors, such as macroeconomic conditions or policy changes, significantly alter the patterns of financial data, rendering the learned model unsuitable for practical implementation. However, a limited time period results in a scarcity of data, diminishing the performance of the algorithm. With our approach, this issue is addressed by generating a sufficient quantity of data from $\mathbb{Q}$, which closely approximates the current real model, $\mathbb{P}$. Additionally, we can retrain $\mathbb{Q}$ with certain schedule to enable long-time implementation of our algorithm (see an example of retraining trick in Subsection 4.3.2.
@@
Initialization: Initialize Q-network $Q_{\alpha}$ and policy network $\pi_{\beta}$; initialize target networks: $\alpha_{\mathrm{targ}}\leftarrow\alpha$, $\beta_{\mathrm{targ}}\leftarrow\beta$; set up an empty replay buffer $\mathcal{D}$ and an empty scenario pool $\mathcal{S}$, both with size $L$; train the score network $s_{\theta}$ and feature extraction module $R_{\theta}$ using Algorithm 2. Collect context window data $s^{-T:-1}$.
@@
$-\frac{\gamma}{2}Q_{\alpha}\big{(}1,0,h^{1},\pi_{\beta}(1,0,h^{1},c),c\big{)}+c.$
@@ 4.3.2 Real data
For training, we use a same‐resolution UNet which is inspired by Ronneberger et al. (2015) as the score network $s_{\theta}$ and an LSTM Hochreiter and Schmidhuber (1997) as the RNN encoder $R_{\theta}$. To improve the performnance, e employ the refined version of OU process: the variance‐preserving stochastic differential equation (VPSDE):
@@
We set $\gamma=3$ and test the four strategies comprehensively using the following metrics: annualized return, annualized volatility, Sharpe ratio, Sortino ratio, maximal drawdown and Calmer ratio. See Table 2 for the performance. We also report the output portfolio for GenTD3 strategy and the wealth trajectory for all four strategies; see Figures 5. From the output tables and figures, we conclude that two generative-model-based strategies outperform S&P 500 and two baselines (in terms of Sharpe and Sortino ratio), justifying the efficiency of the diffusion model. We emphasize that, as a time series predictor, the diffusion model consists of two parts: the score network $s_{\theta}$ and the feature extraction module $R_{\theta}$. They are both trained in an adaptive manner via the proposed Algorithm 2. For the GenMarkowitz strategy, while the feature variables $h^{t}$ are updated online, it treats the portfolio selection as a one-period problem and solves static Markowitz problems for every $t$. This is in contrast to GenTD3 strategy, which deal with a dynamic decision problem. We observe from Table 2 that the GenTD3 strategy outperforms GenMarkowitz in 5 out of 6 metrics, except annualized return. These results support our proposed adaptive sampling method (Algorithm 3), which includes intertemporal structure in predictions. The GenTD3 agent then exploits such intertemporal structure through an reinforcement learning framework (Algorithm 4).
@@
Because it is difficult to specify a level of risk aversion in reality, we need to justify the robustness of our algorithm under the change of $\gamma$. We report the results in Tables 3-4 for a lower risk aversion $\gamma=0.5$ (which is the same as the synthetic experiment) and a higher risk aversion $\gamma=5$. We observe that, although GenMarkwotiz performs exceptional for $\gamma=5$, it crashes for smaller $\gamma=0.5$, even underperforming EW and S&P 500. GenTD3, in contrast, is stable in varying risk aversion and achieves the best results for $\gamma=0.5$. It also has a performance comparable to GenMarkowitz for $\gamma=5$, better than the benchmarks. This sensitivity analysis indicates the importance of the proposed adaptive sampling scheme. We also conclude that, the two strategies based on diffusion model are suitable for different risk aversions. In particular, choosing the appropriate risk aversion (which is not an easy task, though), GenMarkowitz may also have satisfactory performance, justifying the applicability of diffusion models in financial data.
@@ Appendix A On Assumption 3
If $G_{\theta}$ can be chosen such that it is Lipschitz continuous and bounded, then Assumption 3 is satisfied. Indeed, noting that $h_{2}(\tau)<1$, we have
@@
where $M$ is the upper bound of $G_{\theta}$.
In fact, we can construct a bounded supported $G_{\theta}$ which approximate $G$ with any accuracy. To prove this, we need an improved version of universal approximation property of the ReLU network from Chen et al. (2022):
@@
For any $\varepsilon>0$, there exists a ReLU network architecture $\hat{f}$ on $\mathbb{R}^{m}$ such that for any 1-Lipschitz function $f$ defined on $[0,1]^{m}$ with $f(0)=0$, we have $\|f-\hat{f}\|_{\infty}\leq\varepsilon$. Moreover, $\hat{f}$ is $10d$-Lipschitz.
@@
For any $\varepsilon_{\rm score}>0$, there exists a constant $R>0$ and a ReLU network $G_{\theta}$, such that $G_{\theta}(\tau,\cdot,\cdot)$ is supported on $B_{R}:=\{|x^{1:t-1}|\leq R\}\times\{|x|\leq R\}$ for any $\tau\in[\tau_{0},\mathcal{T}]$, and
@@
with a sufficiently large $R$. On the other hand, for some $20<\eta<1/$ 2, if $|x^{t}|\leq\eta|x|$, then $|x-x^{t}h_{1}(\tau)|^{2}\geq\frac{1}{2}|x|^{2}-|h_{1}(\tau)x^{t}|^{2}\geq(\frac{1}{2}-\eta)|x|^{2}$, hence
@@
Combining (A.25) and (A.28) and taking $R$ sufficiently large, we have
@@
From Assumption 2-1 and Lemma A.1333By an additional ReLu layer, the support set can be transferred from $[0,1]^{d}$ to $B_{R}$. See Appendix B.1 of Chen et al. (2023a) for details., by choosing appropriate $\theta$, it can be achieved that
@@
To obtain the Wasserstein bounds from total variation bounds, we first provide a uniform-in-time moment estimate for SDE under strong dissipative condition; see Section 1.2 of Khasminskii (2012). Suppose that for the function $f:[0,\mathcal{T}]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, there exists $C,R_{0}>0$ such that for any $|x|>R_{0}$,
$x\cdot f(t,x)\leq-(1+\delta)|x|^{2}+C.$ (B.1)
@@
$\mathrm{d}X_{t}=(X_{t}+f(t,X_{t}))\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t}.,$ (B.2)
@@
There exist constants $c_{1},c_{2}>0$, which are independent from $t$, such that
$\sup_{t\geq 0}\mathbb{E}[e^{c_{1}|X_{t}|^{2}}]\leq c_{2}.$ (B.4)
@@
Let $V(x)=e^{\theta x}$, where $\theta>0$ will be determined later. Applying Itô’s formula to $V(X_{s})$, we have
[…]
By considering a sequence of localization stopping time $\tau_{n}\to\infty$,
[…]
Choose $\theta<\delta/2$, $R^{2}>2\theta(C+d)/(\theta\delta-2\theta^{2})$ and denote $I_{s}=(4\theta^{2}-2\theta\delta)|X_{s}|^{2}+2\theta(C+d)\big{)}V(X_{s})$. We note that $I_{s}\leq-\gamma V(X_{s})$ with $\gamma:=2\theta(C+d)$ when $|X_{s}|\geq R$, and $I_{s}\leq K$ when $|X_{s}|\leq R$, where $K$ is some constant depending on $\theta,\delta,C,d,R$. We thus obtain
$\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|>R}]\leq$ $\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})]$ (B.8)
$\leq$ $\mathbb{E}[V(X_{0})]-\gamma\int_{0}^{t\wedge\tau_{n}}V(X_{s})I_{|X_{s}|>R}+Kt$ (B.9)
$=$ $\mathbb{E}[V(X_{0})]-\gamma\int_{0}^{t}V(X_{s\wedge\tau_{n}})I_{|X_{s\wedge\tau_{n}}|>R}+Kt.$ (B.10)
@@
$\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|>R}]\leq e^{-\gamma t}\mathbb{E}[V(X_{0})I_{|X_{0}|>R}]+\frac{K}{\gamma}\leq C,$
where the constant $C$ is independent of $t$. On the other hand, clearly $\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|\leq R}]\leq C$, thus by Fatou’s Lemma,
$\mathbb{E}[V(X_{t})]]\leq\liminf_{n\to\infty}\mathbb{E}[V(X_{t\wedge\tau_{n}})]\leq C.$
Thus, (B.4) is established with $c_{1}=\theta$, $c_{2}=C$, a generic constant that is independent of $t$. ∎
@@
$\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})^{2}\leq$ $C\bigg{(}e^{-2\mathcal{T}}+\int_{0}^{\mathcal{T}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|x^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,y^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|x^{1:t})\big{|}^{2}\mathrm{d}\tau\bigg{)}$ (B.14)
@@
$\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})\leq C\big{(}e^{-\mathcal{T}}+\mathcal{E}(x^{1:t})^{1/2}+\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|\big{)}.$ (B.16)
@@
$\mathcal{W}_{2}^{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})\leq C\bigg{(}R^{2}\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})+\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}|X^{t+1}|^{2}I_{\{|X^{t+1}|\geq R\}}\big{]}+\mathbb{E}_{\mathbb{Q}_{y^{1:t}}}\big{[}|Y^{t+1}|^{2}I_{\{|Y^{t+1}|\geq R\}}\big{]}\bigg{)}.$ (B.18)
@@
$\mathcal{W}_{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})^{2}\leq$ $C(R^{2}e^{-\mathcal{T}}+R^{2}\mathcal{E}(x^{1:t})^{1/2}+R^{2}\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|+Ce^{-\frac{c}{2}R}+c_{2}e^{-c_{1}R^{2}})$ (B.25)
$\leq$ $C(R^{2}e^{-\mathcal{T}}+R^{2}\mathcal{E}(x^{1:t})^{1/2}+R^{2}\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|+e^{-cR}).$ (B.26)
@@
(2). Because $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ if and only if $\pi_{x^{1:t},y^{1:t}}\in\Pi(\mathbb{P}_{x^{1:t}},\mathbb{Q}_{y^{1:t}})$ (Backhoff et al. (2017), Proposition 5.1), we construct a specific coupling $\pi^{\varepsilon}\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ as follows:
@@
    $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x^{1}-y^{1}|^{2}\pi^{\varepsilon}_{1}(\mathrm{d}x^{1},\mathrm{d}y^{1})\leq\mathcal{W}_{2}^{2}(\mathbb{P}^{1},\mathbb{Q}^{1})+\varepsilon$;
@@
$\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q})\leq$ $\mathbb{E}_{\pi^{\varepsilon}}|X^{1:T}-Y^{1:T}|^{2}$ (B.41)

### v3

@@ 4.2 Solving portfolio choice problem with $\mathbb{Q}$
For reinforcement learning in financial applications, it is advisable to refrain from utilizing long-term data, as certain global factors, such as macroeconomic conditions or policy changes, significantly alter the patterns of financial data, rendering the learned model unsuitable for practical implementation. However, a limited time period results in a scarcity of data, diminishing the performance of the algorithm. With our approach, this issue is addressed by generating a sufficient quantity of data from $\mathbb{Q}$, which closely approximates the current real model, $\mathbb{P}$. Additionally, we can retrain $\mathbb{Q}$ with certain schedule to enable long-time implementation of our algorithm (see an example of retraining trick in Subsection 4.3.2).
@@
Initialization: Initialize Q-network $Q_{\alpha}$ and policy network $\pi_{\beta}$; initialize target networks: $\alpha_{\mathrm{targ}}\leftarrow\alpha$, $\beta_{\mathrm{targ}}\leftarrow\beta$; set up an empty replay buffer $\mathcal{D}$ and an empty scenario pool $\mathcal{S}$, both with size $L$; train the score network $s_{\theta}$ and RNN encoder $R_{\theta}$ using Algorithm 2. Collect context window data $s^{-T:-1}$.
@@
$\frac{\gamma}{2}Q_{\alpha}\big{(}1,0,h^{1},\pi_{\beta}(1,0,h^{1},c),c\big{)}+c.$
@@ 4.3.2 Real data
For training, we use a same‐resolution UNet which is inspired by Ronneberger et al. (2015) as the score network $s_{\theta}$ and an LSTM Hochreiter and Schmidhuber (1997) as the RNN encoder $R_{\theta}$. To improve the performnance, we employ the refined version of OU process: the variance‐preserving stochastic differential equation (VPSDE):
@@
We set $\gamma=3$ and test the four strategies comprehensively using the following metrics: annualized return, annualized volatility, Sharpe ratio, Sortino ratio, maximal drawdown and Calmer ratio. See Table 2 for the performance. We also report the output portfolio for GenTD3 strategy and the wealth trajectory for all four strategies; see Figures 5. From the output tables and figures, we conclude that two generative-model-based strategies outperform S&P 500 and two baselines (in terms of Sharpe and Sortino ratio), justifying the efficiency of the diffusion model. We emphasize that, as a time-series predictor, the diffusion model consists of two parts: the score network $s_{\theta}$ and the RNN encoder $R_{\theta}$. They are both trained in an adaptive manner via the proposed Algorithm 2. For the GenMarkowitz strategy, while the feature variables $h^{t}$ are updated online, it treats the portfolio selection as a one-period problem and solves static Markowitz problems for every $t$. This is in contrast to GenTD3 strategy, which deal with a dynamic decision problem. We observe from Table 2 that the GenTD3 strategy outperforms GenMarkowitz in 5 out of 6 metrics, except annualized return. These results support our proposed adaptive sampling method (Algorithm 3), which includes intertemporal structure in predictions. The GenTD3 agent then exploits such intertemporal structure through an reinforcement learning framework (Algorithm 4).
@@
Because it is difficult to specify a level of risk aversion in reality, we need to justify the robustness of our algorithm under the change of $\gamma$. We report the results in Tables 3-4 for a lower risk aversion $\gamma=0.5$ and a higher risk aversion $\gamma=5$. We observe that, although GenMarkwotiz performs exceptional for $\gamma=5$, it crashes for smaller $\gamma=0.5$, even underperforming EW and S&P 500. GenTD3, in contrast, is stable in varying risk aversion and achieves the best results for $\gamma=0.5$. It also has a performance comparable to GenMarkowitz for $\gamma=5$, better than the benchmarks. This sensitivity analysis indicates the importance of the proposed adaptive sampling scheme. We also conclude that, the two strategies based on diffusion model are suitable for different risk aversions. In particular, choosing the appropriate risk aversion (which is not an easy task, though), GenMarkowitz may also have satisfactory performance, justifying the applicability of diffusion models in financial data.
@@ Appendix A On Assumption 3
If $G_{\theta}$ can be chosen such that it is Lipschitz continuous and compactly supported in $\{|x^{1:t}|<R\}$, then Assumption 3 is satisfied with $M_{\rm disp}\sim R^{2}$. Indeed, noting that $h_{2}(\tau)<1$, we have
@@
where $M$ is the upper bound of $G_{\theta}$. Because the real score function, hence $G$, is assumed to be Lipchitz (say, with a Lipchitz constant $L$), we have $M\leq\|G-G_{\theta}\|_{\infty}+\sup_{|x^{1:t}|<R}|G(\tau,x^{1:t})|\leq 1+LR$. This is exactly Assumption 3 with $M_{\rm disp}\sim R^{2}$.
In fact, we can indeed construct a bounded supported $G_{\theta}$ which approximate $G$ with any accuracy $\varepsilon_{\rm score}$, with the support size $R$ satisfies $\varepsilon_{\rm score}\sim e^{-cR^{2}}$. To prove this, we need an improved version of universal approximation property of the ReLU network from Chen et al. (2022):
@@
For any $\varepsilon>0$, there exists a ReLU network architecture $\hat{f}$ on $\mathbb{R}^{m}$ such that for any 1-Lipschitz function $f$ defined on $0,1]^{m}$ with $f(0)=0$, we have $\|f-\hat{f}\|_{\infty}\leq\varepsilon$. Moreover, $\hat{f}$ is $10d$-Lipschitz.
@@
For any $\varepsilon_{\rm score}>0$, there exists a constant $R>0$ and a ReLU network $G_{\theta}$, such that $G_{\theta}(\tau,\cdot,\cdot)$ is supported on $B_{R}:=\{|x^{1:t-1}|\leq R\}\times\{|x|\leq R\}$ for any $\tau\in[\tau_{0},\mathcal{T}]$, where $e^{-cR^{2}}\sim\varepsilon_{\rm score}$ and
@@
with a sufficiently large $R$. On the other hand, for some $0<\eta<1/2$, if $|x^{t}|\leq\eta|x|$, then $|x-x^{t}h_{1}(\tau)|^{2}\geq\frac{1}{2}|x|^{2}-|h_{1}(\tau)x^{t}|^{2}\geq(\frac{1}{2}-\eta)|x|^{2}$, hence
@@
Combining (A.25) and (A.28) and taking $R$ sufficiently large in the order $e^{-cR^{2}}\sim\varepsilon_{\rm score}$, we have
@@
From Assumption 2-1 and Lemma A.1333By an additional ReLu layer, the support set can be transferred from $0,1]^{d}$ to $B_{R}$. See Appendix B.1 of Chen et al. (2023a) for details., by choosing appropriate $\theta$, it can be achieved that
@@
To obtain the Wasserstein bounds from total variation bounds, we first provide a uniform-in-time moment estimate for SDE under strong dissipative condition; see Section 1.2 of Khasminskii (2012). Suppose that for the function $f:[0,\mathcal{T}]\times\mathbb{R}^{d}\to\mathbb{R}^{d}$, there exists $M_{\rm disp},\delta>0$ such that for any $t,x$,
$x\cdot f(t,x)\leq-(1+\delta)|x|^{2}+M_{\rm disp}.$ (B.1)
@@
$\mathrm{d}X_{t}=(X_{t}+f(t,X_{t}))\mathrm{d}t+\sqrt{2}\mathrm{d}B_{t},$ (B.2)
@@
There exist constants $c_{1},c_{2},c_{3}>0$, which are independent from $t$ and $M_{\rm disp}$, such that
$\sup_{t\geq 0}\mathbb{E}[e^{c_{1}|X_{t}|^{2}}]\leq c_{2}e^{c_{3}M_{\rm disp}}.$ (B.4)
@@
Let $V(x)=e^{\theta|x|^{2}}$, where $\theta>0$ will be determined later. Applying Itô’s formula to $V(X_{s})$, we have
$V(X_{t})-V(X_{0})=$ $\int_{0}^{t}\bigg{(}2\theta\Big{(}V(X_{s})X_{s}\cdot\big{(}X_{s}+f(s,X_{s})\big{)}\Big{)}+2\theta dV(X_{s})+4\theta^{2}|X_{s}|^{2}V(X_{s})\bigg{)}ds$ (B.5)
$+2\sqrt{2}\theta\int_{0}^{t}X_{s}V(X_{s})dB_{s}.$ (B.6)
Considering a sequence of localization stopping time $\tau_{n}\to\infty$ and taking expectation on the above identity (B.5), we obtain
$\mathbb{E}[V(X_{t\wedge\tau_{n}})]-\mathbb{E}[V(X_{0})]=$ $\int_{0}^{t\wedge\tau_{n}}\bigg{(}2\theta\Big{(}\mathbb{E}\big{[}V(X_{s})X_{s}\cdot\big{(}X_{s}+f(s,X_{s})\big{)}\big{]}\Big{)}$ (B.7)
$+2\theta d\mathbb{E}[V(X_{s})]+4\theta^{2}\mathbb{E}[|X_{s}|^{2}V(X_{s})]\bigg{)}ds$ (B.8)
$=$ $\int_{0}^{t}\bigg{(}2\theta\Big{(}\mathbb{E}\big{[}V(X_{s\wedge\tau_{n}})X_{s\wedge\tau_{n}}\cdot\big{(}X_{s\wedge\tau_{n}}+f(s\wedge\tau_{n},X_{s\wedge\tau_{n}})\big{)}\big{]}\Big{)}$ (B.9)
$+2\theta d\mathbb{E}[V(X_{s\wedge\tau_{n}})]+4\theta^{2}\mathbb{E}[|X_{s\wedge\tau_{n}}|^{2}V(X_{s\wedge\tau_{n}})]\bigg{)}ds$ (B.10)
The differential form yields
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}[V(X_{t\wedge\tau_{n}})]=$ $2\theta\mathbb{E}\big{[}V(X_{s\wedge\tau_{n}})X_{s\wedge\tau_{n}}\cdot\big{(}X_{s\wedge\tau_{n}}+f(s\wedge\tau_{n},X_{s\wedge\tau_{n}})\big{)}\big{]}\Big{)}$ (B.11)
$+2\theta d\mathbb{E}[V(X_{t\wedge\tau_{n}})]+4\theta^{2}\mathbb{E}[|X_{t\wedge\tau_{n}}|^{2}V(X_{t\wedge\tau_{n}})]$ (B.12)
$\leq$ $(4\theta^{2}-2\theta\delta)\mathbb{E}[|X_{t\wedge\tau_{n}}|^{2}V(X_{t\wedge\tau_{n}})]+2\theta(M_{\rm disp}+d)\mathbb{E}[V(X_{t\wedge\tau_{n}})].$ (B.13)
Choose $\theta<\delta/2$, $R^{2}>2(M_{\rm disp}+d)/(\delta-2\theta)$ and denote $I_{t}=\big{(}(4\theta^{2}-2\theta\delta)|X_{t\wedge\tau_{n}}|^{2}+2\theta(M_{\rm disp}+d)\big{)}V(X_{t\wedge\tau_{n}})$. On the one hand, with $\gamma:=2\theta(M_{\rm disp}+d)$, it is clear that $I_{t}\leq-\gamma V(X_{t\wedge\tau_{n}})$ when $|X_{t\wedge\tau_{n}}|\geq R$. On the other hand, when $|X_{t\wedge\tau_{n}}|\leq R$,
$I_{t}\leq$ $2\theta(M_{\rm disp}+d)V(X_{t\wedge\tau_{n}})$ (B.14)
$\leq$ $2\theta(M_{\rm disp}+d)V(R)$ (B.15)
$=$ $2\theta(M_{\rm disp}+d)e^{\frac{2\theta(M_{\rm disp}+d)}{\delta-2\theta}}$ (B.16)
$\leq$ $c^{\prime}_{1}e^{c^{\prime}_{2}M_{\rm disp}},$ (B.17)
where $c^{\prime}_{1}$, $c^{\prime}_{2}$ are some constants independent of $t$ and $M_{\rm disp}$. We thus obtain
$\frac{\mathrm{d}}{\mathrm{d}t}\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})]\leq$ $-\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|>R}]+c^{\prime}_{1}e^{c^{\prime}_{2}M_{\rm disp}}$ (B.18)
$=$ $-\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})]+\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})I_{|X_{t\wedge\tau_{n}}|\leq R}]+c^{\prime}_{1}e^{c^{\prime}_{2}M_{\rm disp}}$ (B.19)
$\leq$ $-\gamma\mathbb{E}[V(X_{t\wedge\tau_{n}})]+c^{\prime}_{3}e^{c^{\prime}_{4}M_{\rm disp}}$ (B.20)
@@
$\mathbb{E}\big{[}V(X_{t\wedge\tau_{n}})]\leq e^{-\gamma t}\mathbb{E}[V(X_{0})I_{|X_{0}|>R}]+\frac{c^{\prime}_{3}e^{c^{\prime}_{4}M_{\rm disp}}}{\gamma}\leq c^{\prime}_{5}e^{c_{6}^{\prime}M_{\rm disp}},$
Here, constants $c^{\prime}_{1}\sim c^{\prime}_{6}$ are all independent of $t$, $n$ and $M_{\rm disp}$. Thus, letting $n\to\infty$, (B.4) is established with $c_{1}=\theta$, $c_{2}=c^{\prime}_{5}$, and $c_{3}=c^{\prime}_{6}$, all independent of $t$ and $M_{\rm disp}$. ∎
@@
$\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})^{2}\leq$ $C\bigg{(}e^{-2\mathcal{T}}+\int_{0}^{\mathcal{T}}\mathbb{E}_{X_{\tau}^{t+1}\sim p_{t+1}(\tau,\cdot|x^{1:t})}\big{|}s^{t+1}_{\theta}(\tau,y^{1:t},X_{\tau}^{t+1})-\nabla_{x}\log p_{t+1}(\tau,X_{\tau}^{t+1}|x^{1:t})\big{|}^{2}\mathrm{d}\tau\bigg{)}$ (B.24)
@@
$\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})\leq C\big{(}e^{-\mathcal{T}}+\mathcal{E}(x^{1:t})^{1/2}+\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|\big{)}.$ (B.26)
@@
$\mathcal{W}_{2}^{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})\leq C\bigg{(}R^{2}\mathrm{TV}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})+\mathbb{E}_{\mathbb{P}_{x^{1:t}}}\big{[}|X^{t+1}|^{2}I_{\{|X^{t+1}|\geq R\}}\big{]}+\mathbb{E}_{\mathbb{Q}^{\mathcal{T}}_{y^{1:t}}}\big{[}|Y^{t+1}|^{2}I_{\{|Y^{t+1}|\geq R\}}\big{]}\bigg{)}.$ (B.28)
@@
$\mathcal{W}_{2}(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})^{2}\leq$ $C(R^{2}e^{-\mathcal{T}}+R^{2}\mathcal{E}(x^{1:t})^{1/2}+R^{2}\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|+Ce^{-\frac{c}{2}R}+e^{-c_{1}R^{2}+c_{2}M_{\rm disp}})$ (B.35)
$\leq$ $C(R^{2}e^{-\mathcal{T}}+R^{2}\mathcal{E}(x^{1:t})^{1/2}+R^{2}\mathcal{T}^{1/2}|x^{1:t}-y^{1:t}|+e^{-cR+cM_{\rm disp}}).$ (B.36)
@@
(2). Because $\pi\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})$ if and only if $\pi_{x^{1:t},y^{1:t}}\in\Pi(\mathbb{P}_{x^{1:t}},\mathbb{Q}^{\mathcal{T}}_{y^{1:t}})$ (Backhoff et al. (2017), Proposition 5.1), we construct a specific coupling $\pi^{\varepsilon}\in\Pi_{\rm bc}(\mathbb{P},\mathbb{Q})$ as follows:
@@
    $\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}|x^{1}-y^{1}|^{2}\pi^{\varepsilon}_{1}(\mathrm{d}x^{1},\mathrm{d}y^{1})\leq\mathcal{W}_{2}^{2}(\mathbb{P}_{1},\mathbb{Q}^{\mathcal{T}}_{1})+\varepsilon$;
@@
$\mathcal{A}\mathcal{W}_{2}^{2}(\mathbb{P},\mathbb{Q}^{\mathcal{T}})\leq$ $\mathbb{E}_{\pi^{\varepsilon}}|X^{1:T}-Y^{1:T}|^{2}$ (B.51)